OpenStudy (anonymous):

If the length of an arc is 12pi inches and the radius of the circle is 10 inches, what is the measure of the arc?

OpenStudy (anonymous):

θ = s/r = 12pi/10 = 3.77 radians measure = radians * 180/pi = 3.77*(180/22/7) = 216 degree

OpenStudy (anonymous):

hey ive got that wrong it should be length of arc = a * c so a = l / c = 12pi/10 and soon as i saw you typing i realised my mistake math - thanx

OpenStudy (anonymous):

sorry cherri

OpenStudy (anonymous):

always welcome